The Network Crew says the AP (the device that transmits and receives data for this location) has failed and will be replaced tomorrow morning likely before 9 a.m. (Wednesday, January 21st, 2009).
An update will be forth coming.
[2009-01-21 11:44:45 | Philip Loarie]
Networks reports the problem persists however, they are currently working on the problem and hope to have a solution within the next 2 hours (about 1:30 p.m. on Jan. 21st, 2009).
[2009-01-21 16:14:37 | Philip Loarie]
The BID center wireless outage has been resolved by the EECS IRIS Network staff today. They report that the issue was resolved earlier today but it was not clear due to their monitoring software not being configured correctly.
Resolved as of 2009-01-21 16:09:00